# An All-digital 65-nm Tsetlin Machine Image Classification Accelerator with 8.6 nJ per MNIST Frame at 60.3k Frames per Second

### v1

@@ IV Design Overview
The block diagram of the ConvCoTM accelerator is shown in Fig. 2. The accelerator includes all circuitry required for performing inference, and only minimal interaction with the system processor is required. The three main modules are i) a data interface for transfer of model parameters and image samples, ii) a module for model storage and iii) an inference core that performs classification of an image.
Figure 2: ConvCoTM accelerator ASIC block diagram.
The main configuration parameter of a TM solution is the number of clauses. Due to chip area restrictions, we decided to implement a configuration with 128 clauses. The following subsections describe the different sub-modules and the accelerator’s operation in detail.
@@ IV-A Interface to the System Processor
The data interface towards the external system processor, is an 8-bit parallel type inspired by the ARM AXI Stream interface [50]. It is used for transferring model data to the accelerator during load model mode, and for transferring images in inference mode. When a classification operation is finished, the accelerator outputs an interrupt. In addition, it outputs a separate 8-bit signal consisting of the predicted class (4 bits) together with the true label of the sample (4 bits).
@@ IV-C Image Data Buffer and Patch Generation
A data buffer is included with room for two complete $28\times 28$ booleanized images. While one image is being processed, another image can be transferred from the external system processor to the other image buffer, thereby improving throughput. We denote this as the continuous classification mode. A complete booleanized image requires $28\times 28/8=98$ bytes. In addition, one byte is allocated for the image label (0 to 9).
@@
During the convolution, the window is moved to the right with increments of one. As described in subsection III-C, the position of the window is thermometer encoded. There are 19 different $x$-coordinates for the window, requiring 18 bits. When the window has been evaluated at the rightmost $x$-position (0,18), the contents of all 10 rows are shifted upwards with one step, and the next image datarow is loaded into the lowermost row. Then the window starts again at the leftmost position, now with patch coordinates (1,0), and moves to the right in steps of one. This procedure is repeated until all 28 datarows of the image have been processed. Also for the $x$-coordinate, there are 19 different positions. During a complete convolution, $19\times 19=361$ patches are generated.
@@ IV-D Clause Pool
The ConvCoTM is configured with 128 clauses that operate in parallel. Each of the 128 clauses takes the literals, $[l_{0},l_{1},\dots,l_{271}]$, as input, as well as their corresponding TA action signals, $[i_{0},i_{1},\dots,i_{271}]$, from the model register. Fig. 3 shows a simplified circuit diagram of the clause logic. A literal will be included in the logical AND expression, that constitutes the clause, if its corresponding TA action signal is high. A special situation occurs when a clause is empty, i.e., when there are no include TA actions. In this case, additional logic will set the signal Empty high, forcing $c_{j}^{b}$ low.
The DFF in Fig. 3 constitutes a single-bit register for clause $j$. During the convolution, its content, $c_{j}$, is ORed with the combinational clause value, $c_{j}^{b}$, of the current patch, thus implementing the sequential OR function in Eq. (6). Before starting a new convolution (patch generation), the DFF is reset.
The signal $c_{j}$ is also fed back to the input of the OR-gates, as there is no need for the combinational logic to evaluate a clause for more patches if $c_{j}$ already is high. This feedback mechanism, hereinafter denoted the clause switching reduction feedback (CSRF), reduces the digital switching in the combinational part of the clause logic. Simulations on MNIST image classification with CSRF enabled, showed an average of 50% reduction in the toggling rate of $c_{j}^{b}$ for each clause. The accelerator ASIC has a dedicated pin that can enable or disable the CSRF.
@@ IV-E Inference
The inference operation starts with loading an image into the data buffer, followed by a convolution process, as described in Subsection IV-C. When finished, each clause value, $c_{j}$, is weighted, for each class, according to Eq. (3).
All ten class sums are generated in parallel, by adders configured in a tree structure and utilizing a three-stage pipeline. Finally, an argmax module [18] selects the class with the highest class sum, as the predicted class, $\hat{y}$. Fig. 4 shows a simplified state diagram of the accelerator.
@@
The accelerator’s latency, measured from when the system processor initiates the first byte transfer of the image, until a prediction result is available, is 471 clock cycles. This includes 99 clock cycles for transferring the 98 image bytes and the label byte, and 372 clock cycles for the patch generation and class prediction.
The accelerator can also operate in continuous mode, where a new sample is loaded into the image buffer while the current sample is being processed, see Subsection IV-C. This increases the throughput, as image samples can be processed every 372’th clock cycle. Any timing overhead in the system processor will add to the total latency.
@@ IV-F Clock Domains
The key to reducing the digital switching power of the accelerator ASIC lies at the top level of the ASIC architecture. Two separate clock domains are utilized, each with its own dedicated clock pin, assigned to i) the model part and ii) the inference core.
When a model has been transferred from the system processor to the ASIC, the TA action signals and the weights are available in registers which do not need further clocking. By stopping the clock to the model part when operating in inference mode, the digital switching power is reduced significantly, as the model part constitutes approximately 90% of the accelerator’s DFFs. To further reduce power consumption, standard clock gating is applied for the inference module. This ensures, for instance, that the pipelined registers associated with the class sum generation are enabled and clocked only for 4 clock cycles per classification cycle. The clock-gating can be enabled/disabled by an external pin.
@@ V Implementation and Measurement Results
The ConvCoTM accelerator was designed in VHDL and first implemented and verified on a Xilinx ZCU104 FPGA development board. The ASIC was then implemented in a 65 nm low-leakage CMOS technology from UMC, and packaged in a JLCC44 ceramic package. Fig. 5 shows a photograph and a layout plot of the ASIC. The accelerator’s core area is 2.7 mm2.
Figure 5: Photo and layout plot of the ConvCoTM accelerator chip.
A test PCB was developed for the ASIC. This board was connected to a Digilent Zybo Z7-20 development board equipped with a Xilinx Zynq XC7Z020 FPGA, as shown in Fig. 6. One of the FPGA’s ARM9 cores was configured as system processor, and the FPGA board provided clocks, model data, image data samples and control signals to the ASIC accelerator. The processor also read back accelerator status/interrupt signals and the predicted/true class per sample. To interface with the FPGA board, the ASIC’s input/output (IO) ports operated at 3.3 V supply voltage. The accelerator core operated from 0.82 V to 1.2 V.
@@
For each of the datasets MNIST, FMNIST and KMNIST, the Tsetlin Machine Unified (TMU) SW-version [51] of the ConvCoTM was trained to find suitable models. Maximum/minimum limits were set on the clause weights to fit with the allocated 8-bits per weight in the accelerator. For simplicity, the images from the test dataset were preprocessed (booleanized), as described in subsection III-D, and included in the FPGA processor’s test program.
Model data and image samples were applied from the FPGA to the accelerator via the AXI Stream interface. We utilized DMA functionality on the FPGA to enable fast data transfer. The maximum classification rate achieved, including system processor timing overhead, was 60.3k images per second in continuous mode, with an accelerator clock frequency of 27.8MHz. The speed limiting factor was the electrical interface between the FPGA and the ASIC. The latency when classifying a single image was 25.4 $\mu$s, including the transfer time of the image to the accelerator as well as timing overhead in the system processor.
@@
In a real application, the ConvCoTM design would typically be included as a processor peripheral module in an SoC. Therefore, we considered the power consumption of the accelerator core as the most interesting performance power parameter, and our measurements excluded energy consumed by the digital 3.3 V IOs of the ASIC. Table I summarizes the characteristics and performance of the accelerator.
@@ 21 $\mu$W b,d
Latency (single image classi-, 25.4 $\mu$s a
@@
For power consumption measurements, a Joulescope JS220 precision energy analyzer was used. The power consumption during inference was measured while the accelerator was set in a test mode with repeated classifications of the full 10k sample dataset. The power consumed by the accelerator core with 1.20 V supply voltage and a clock frequency of 27.8 MHz was 1.15 mW. This corresponds to an EPC of 19.1 nJ. With a supply voltage of 0.82 V, the power consumption was 0.52 mW and the corresponding EPC was 8.6 nJ. With 1.2 V core supply voltage and a clock frequency of 1 MHz, the achieved classification rate was 2.27k images/second at a power consumption of 81 $\mu$W.
@@ VI Discussions
We consider the ConvCoTM solution with 128 clauses as a small configuration. In light of this, we are satisfied with the obtained test accuracy which is comparable with other low-complexity and low-power solutions. Table II shows the performance and properties of state-of-the-art low-power IC implementations for image classification. Inspired by the TinyML [52] benchmarks for embedded systems we specify the classification rate, the peak power consumption and the EPC. These are system parameters which are independent of the accelerator model and architecture [53]. However, when used for comparing solutions, one must be careful to specify the ML task performed. The ICs shown in Table II operate on images and datasets that range from low to high complexity. Accelerators that operate on CIFAR-10 and other more complex images, typically have lower classification rates and higher EPC compared to accelerators that operate on $28\times 28$ images.
@@
This work TCAS-II’23 Yejun [38] JSSC’23       Yang [9] TCAS-I’20 Mauro [6] JSSC’21       Knag [7] TCAS-I’20 Bankman [5] JSSC’17 Chen [35] JETCAS’19 Chen [37]
Technology 65 nm CMOS 65 nm CMOS 40 nm CMOS 22 nm FD-SOI 10 nm FinFET 28 nm CMOS 65 nm CMOS 65 nm CMOS
Active area 2.7 mm2 0.57 mm2 0.98 mm2 2.3 mm2 0.39 mm2 4.6 mm2 12.3 mm2 Not stated
Algorithm ConvCoTM SNN Ternary CNN BNN BNN BNN CNN Sparse DNN
@@
Design type Digital Neuromorphic mixed-signal IMC, mixed-signal Digital (SoC) Digital IMC, mixed-signal Digital Digital
Image dataset MNIST, FMNIST, KMNIST MNIST MNIST CIFAR-10 CIFAR-10 CIFAR-10 Imagenet Imagenet
Measured test accuracy 97.42%, 84.54%, 82.55% 95.35% 97.1% 99% of nominal accuracy 86% 86% Not stated 79.6% i           79.7% j
Classifications per second 60.3k a           2.27k b 233k (1.2 V)           40k (0.7 V) 549 15.4 Not stated 237 35 g           0.7 h 278 f,i           1470 f,j
Latency e 25.4 $\mu$s a           0.66 ms b Not stated Not stated Not stated Not stated Not stated 115.3 ms Not stated
Inference power 1.15 mW a,c 0.52 mW a,d 81 $\mu$b,c 21 $\mu$b,d 9.37 mW (1.2 V) 0.517 mW (0.7 V) 96 $\mu$W 674 $\mu$W 5.6 mW (13 MHz) 607 mW (622 MHz) 0.9 mW 278 mW g 236 mW h Not stated
Energy per classification (EPC) 19.1 nJ a,c           8.6 nJ a,d           35.3 nJ b,c           9.6 nJ b,d 40.17 nJ (1.2 V) 12.92 nJ (0.7 V) 0.18 $\mu$J 43.8 $\mu$J Not stated 3.8 $\mu$J 7.9 mJ g           337 mJ h 1.5 mJ f,i 0.4 mJ f,j
@@
a 27.8 MHz clock frequency. b 1.0 MHz clock frequency. c The accelerator core supply voltage (vdd) is 1.20 V. d vdd=0.82 V.e For processing of a single image sample, including data transfer. f Simulated results. g AlexNet. h VGG-16. i Sparse AlexNet. j Mobilenet.
@@
The mixed-signal IMC-based solution in [9] operates with a very low peak power consumption of 96 $\mu$W and consumes 0.18 $\mu$J per classification on MNIST with a test accuracy of  97.1%. Its classification rate is 549. In [38] a mixed-signal SNN IC accelerator achieves a very low EPC of 12.92 nJ with 0.7 V supply voltage. Operating at 1.2 V the EPC is 40.17 nJ. It utilizes a time-domain echo state network (ESN) [54] approach and thus requires non-standard interfacing with the surrounding system’s resources. It obtains a test accuracy on MNIST of 95.35%.
@@
To the authors’ knowledge, the ConvCoTM accelerator core’s EPC of 8.6 nJ at 0.82 V operating voltage is the lowest reported for a manufactured ASIC that classifies $28\times 28$ greyscale images and which has comparable test accuracy. At 1.2 V operating voltage it obtains an EPC of 19.1 nJ, which is also very low. As the ConvCoTM ASIC is fully digital, the classification rate and thereby also the power consumption, can easily be scaled by adjusting the clock frequency. E.g, with a clock frequency of 1 MHz, the power consumption achieved is only 81 $\mu$W. As the accelerator is all-digital, the test accuracy is independent of the clock frequency.
@@
Several analog/mixed-signal IC accelerators obtain very low power consumption [38, 9, 5]. Such solutions are susceptible to variations in process parameters, supply voltage and operating temperature (PVT), which can affect the performance. All-digital solutions avoid such non-ideal effects [6, 7, 8]. In addition, they are easier to integrate in larger SoCs, and can be scaled and re-targeted to other technologies. This applies for CNNs as well as for TMs.
@@
In the ConvCoTM accelerator, the combinational clause logic draws only a small amount of energy compared to the clock tree of the inference-core DFFs. This is the main reason why the CSRF technique, see Subsection IV-D, did not provide a greater reduction in EPC. Other convolutional TM architectures, that operate at higher clock frequencies, process more literals in parallel, and utilize sparsity in TM models, can potentially benefit more from the CSRF technique.
The ConvCoTM accelerator’s architecture described in this paper is highly suited for scaling. It can be modified to operate on larger images including 3-channel (RGB) images, e.g., based on the TM Composites architecture [31, 32]. From the very low energy consumption obtained for our ConvCoTM accelerator, we envisage that a TM Composites-based solution will perform well compared to previous works.
VII Conclusion

### v2

@@ IV Design Overview
The block diagram of the ConvCoTM accelerator is shown in Fig. 1. The accelerator includes all circuitry required for performing inference, and only minimal interaction with the system processor is required. The three main modules are i) a data interface for transfer of model parameters and image samples, ii) a module for model storage and iii) an inference core that performs classification of an image. The following subsections describe the different sub-modules and the accelerator’s operation in detail.
@@ IV-A Interface to the System Processor
The 8-bit parallel data interface towards the external system processor, is inspired by the ARM AXI Stream interface [36]. It is used for transferring model data to the accelerator during load model mode, and for transferring images in inference mode. When a classification operation is finished, the accelerator outputs an interrupt. In addition, it outputs a separate 8-bit signal consisting of the predicted class of the sample together with the true label (if provided).
@@ IV-C Image Data Buffer and Patch Generation
The accelerator includes a data buffer with room for two complete $28\times 28$ booleanized images. While one image is being processed, another image can be transferred from the external system processor to the other image register, thereby improving throughput. We denote this as continuous classification mode. The size of a complete booleanized image is $28\times 28/8$ bytes $=98$ bytes. One additional byte is allocated for the image label (0 to 9).
@@
During the patch generation, the window is moved to the right with increments of one. As described in Subsection III-C, the position of the window is thermometer encoded. There are 19 different $x$-coordinates for the window, requiring 18 bits. When the window has been evaluated at the rightmost $x$-position ($0,18$), the contents of all 10 rows are shifted upwards with one step, and the next image datarow is loaded into the lowermost register row. The convolution window then starts again at the leftmost position, now with patch coordinates ($1,0$). This procedure is repeated until all 28 datarows of the image have been processed. There are 19 different positions also for the $y$-coordinate, and during a complete patch generation phase, $19\times 19=361$ patches are generated.
@@ IV-D Clause Pool
The ConvCoTM is configured with 128 clauses that operate in parallel. Each clause takes the literals, $[l_{0},l_{1},\dots,l_{271}]$, as input, as well as their corresponding TA action signals, $[i_{j,0},i_{j,1},\dots,i_{j,271}]$, $j\in\{0,\ldots,127\}$, from the model register. Fig. 2 shows a simplified circuit diagram of the clause logic. A literal will be included in the logical AND expression, that constitutes the clause, if its corresponding TA action signal is high. A special situation occurs when a clause is empty, i.e., when there are no include TA actions. In this case, additional logic will set the signal Empty high, forcing $c_{j}^{b}$ low.
The DFF in Fig. 2 constitutes a single-bit register for clause $j$. Its content, $c_{j}$, is ORed with the combinational clause value, $c_{j}^{b}$, of the current patch, during the patch generation. This implements the sequential OR function in Eq. (6). The DFF is reset before starting a new convolution process.
The signal $c_{j}$ is also fed back to the input of the OR-gates, as there is no need for the combinational logic to evaluate a clause for more patches if $c_{j}$ already is high. This feedback, hereinafter denoted the clause switching reduction feedback (CSRF), reduces the switching in the combinational part of the clause logic. Simulations of MNIST classification utilizing CSRF, showed an average of 50% reduction in the toggling rate of $c_{j}^{b}$ for each clause. The accelerator ASIC has a dedicated pin that can enable or disable the CSRF.
@@ IV-E Inference
The inference operation starts with loading an image into the data buffer, followed by sequential generation of the patches, as described in Subsection IV-C. When finished, each clause value, $c_{j}$, is weighted, for each class, according to Eq. (3).
All ten class sums are generated in parallel, by adders configured in a tree-structure and utilizing a three-stage pipeline. Finally, an argmax module [11] selects the class with the highest class sum, as the predicted class, $\hat{y}$. Fig. 3 shows a simplified state diagram of the accelerator.
@@
The accelerator’s latency is 471 clock cycles. This is measured from when the system processor initiates the first byte transfer of the image, until a prediction is available. It includes 99 clock cycles for transferring the 98 image bytes and the label byte, and 372 clock cycles for the patch generation, class summation and class prediction.
When the accelerator operates in continuous mode, a new sample is loaded into the image buffer while the current sample is being processed, see Subsection IV-C. This increases the throughput, as image samples can be processed every 372’th clock cycle. Any timing overhead in the system processor will add to the total latency.
@@ IV-F Clock Domains
The key to reducing the digital switching power of the accelerator, lies at the top level of the ASIC architecture. Two separate clock domains are utilized, each with its own dedicated clock pin, assigned to the model part and the inference core, respectively.
When a model has been transferred from the system processor to the ASIC, the TA action signals and the weights are available in registers that do not need further clocking. By stopping the clock to the model part, when operating in inference mode, the digital switching power is reduced significantly, as the model part constitutes approximately 90% of the accelerator’s DFFs. To further reduce power consumption, standard clock gating is applied for the inference module. This ensures, for instance, that the pipelined registers associated with the class sum generation are enabled and clocked only for four clock cycles per classification phase. The clock-gating can be enabled/disabled by an external pin.
Figure 4: Photo and layout plot of the ConvCoTM accelerator chip.
@@ V Implementation and Measurement Results
The ConvCoTM accelerator was designed in VHDL and implemented in a 65 nm low-leakage CMOS technology from UMC. A JLCC44 ceramic package housed the chip. Fig. 4 shows a photograph and a layout plot of the ASIC. The accelerator’s core area is approximately 2.7 mm2. The VHDL source code for the ASIC implementation (used for initial FPGA verification) is available at a GitHub repository [37].
A printed circuit board (PCB) was developed for testing and characterization of the ASIC. The ASIC PCB was connected to a Digilent Zybo Z7-20 development board, equipped with a Xilinx Zynq XC7Z020 FPGA, as shown in Fig. 5. One of the FPGA’s ARM9 cores was configured as system processor, and the FPGA board provided clocks, model data, image data samples and control signals to the ASIC PCB. The processor also read back accelerator status signals, and the predicted class per sample. To interface with the FPGA board, the ASIC’s input/output (IO) ports operated at 3.3 V supply voltage. The accelerator core operated from 0.82 V to 1.2 V.
@@
For each of the datasets MNIST, FMNIST and KMNIST, the Tsetlin Machine Unified (TMU) SW-version [38] of the ConvCoTM was trained to find suitable models. Maximum/minimum limits were set on the clause weights to fit with the allocated 8-bits per weight in the accelerator. For simplicity, the images from the test datasets were preprocessed (booleanized), as described in Subsection III-D, and included in the FPGA processor’s test program.
Model data and image samples were applied from the FPGA to the accelerator via the AXI Stream interface. We utilized direct memory access (DMA) functionality on the FPGA to enable fast data transfer. The maximum classification rate achieved, including system processor timing overhead, was 60.3k images per second in continuous mode, with an accelerator clock frequency of 27.8MHz. The maximum operating frequency was limited by the general purpose IOs (GPIOs) on the FPGA board. When classifying a single image, the latency was 25.4 $\mu$s, including the transfer time of the image to the accelerator as well as timing overhead in the system processor.
@@ 21 $\mu$W b,d
Latency (single image classi- 25.4 $\mu$s a
@@
For power consumption measurements, a Joulescope JS220 precision energy analyzer was used. The power consumption during inference was measured while the accelerator was set in a test mode with repeated classifications of the full 10k sample dataset. The power consumed by the accelerator core with 1.20 V supply voltage and a clock frequency of 27.8 MHz was 1.15 mW. This corresponds to an EPC of 19.1 nJ. With a supply voltage of 0.82 V, the accelerator core’s power consumption was 0.52 mW and the corresponding EPC was 8.6 nJ. With 1.2 V core supply voltage and a clock frequency of 1 MHz, the achieved classification rate was 2.27k FPS at a power consumption of 81 $\mu$W.
Table I summarizes the characteristics and performance of the accelerator. In a real application, the ConvCoTM design would typically be included as a peripheral module to a microcontroller in an SoC. Therefore, we considered the power consumption of the accelerator core as the most interesting power parameter. In inference mode, the digital 3.3 V IOs, which were required for interfacing to the FPGA board, had a power consumption of 0.76 mW, at a clock frequency of 27.8 MHz.
@@ VI Discussions
We consider the ConvCoTM accelerator with 128 clauses as a small configuration. In light of this, we are satisfied with the obtained test accuracy which is comparable with other low-complexity and ULP solutions. Table II shows the performance and properties of state-of-the-art ULP IC implementations for classification of MNIST and CIFAR-10 datasets. Inspired by the TinyML [39] benchmarks for embedded systems, we specify the classification rate, the peak power consumption and the EPC. These are system parameters which are independent of the accelerator model and architecture [40]. However, when used for comparing solutions, one must be careful to specify the ML task performed. The CIFAR-10 dataset is significantly more challenging than MNIST. Therefore, accelerators that operate on CIFAR-10, have lower classification rates and higher EPC compared to accelerators that operate on MNIST.
@@
This work TCAS-I´25 Zhao [20] TCAS-II’23 Yejun [21] JSSC’23       Yang [9] TCAS-I’20 Mauro [6] JSSC’21       Knag [7] TCAS-I’20 Bankman [5]
Technology 65 nm CMOS 28 nm CMOS 65 nm CMOS 40 nm CMOS 22 nm FD-SOI 10 nm FinFET 28 nm CMOS
Active area 2.7 mm2 0.261 mm2 0.57 mm2 0.98 mm2 2.3 mm2 0.39 mm2 4.6 mm2
Algorithm ConvCoTM CNN SNN Ternary CNN BNN BNN BNN
@@
Design type Digital Analog, time domain Neuromorphic mixed-signal IMC, mixed-signal Digital (SoC) Digital IMC, mixed-signal
Image dataset MNIST, FMNIST, KMNIST MNIST MNIST MNIST CIFAR-10 CIFAR-10 CIFAR-10
Measured test accuracy 97.42%, 84.54%, 82.55% 97.9% 95.35% 97.1% 99% of nominal accuracy 86% 86%
Classifications per second 60.3k a           2.27k b 3508 233k (1.2 V)           40k (0.7 V) 549 15.4 Not stated 237
Latency e 25.4 $\mu$s a           0.66 ms b Not stated Not stated Not stated Not stated Not stated Not stated
Inference power 1.15 mW a,c 0.52 mW a,d 81 $\mu$b,c 21 $\mu$b,d 11.6 $\mu$W 9.37 mW (1.2 V) 0.517 mW (0.7 V) 96 $\mu$W 674 $\mu$f 5.6 mW (13 MHz) 607 mW (622 MHz) 0.9 mW
Energy per classification (EPC) 19.1 nJ a,c           8.6 nJ a,d           35.3 nJ b,c           9.6 nJ b,d 3.32 nJ 40.17 nJ (1.2 V) 12.92 nJ (0.7 V) 0.18 $\mu$J 43.8 $\mu$J Not stated 3.8 $\mu$J
@@
a 27.8 MHz clock frequency. b 1.0 MHz clock frequency. c The accelerator core supply voltage (vdd) is 1.20 V. d vdd=0.82 V.e For processing of a single image sample, including data transfer. f Full SoC.
@@
The mixed-signal IMC-based solution in [9] operates with a peak power consumption of 96 $\mu$W and consumes 0.18 $\mu$J per classification on MNIST, with a test accuracy of  97.1%. In [21] a mixed-signal SNN IC accelerator achieves an EPC of 12.92 nJ with 0.7 V supply voltage. It utilizes a time-domain echo state network (ESN) [41] approach and thus requires non-standard interfacing with the surrounding system’s resources. The test accuracy achieved on MNIST is 95.35%. The 28 nm CMOS CNN-based accelerator in [20] utilizes an IMC architecture with analog and time-domain signal processing. It achieves an EPC of 3.32 nJ, which is the lowest reported for MNIST.
@@
Our ConvCoTM accelerator ASIC achieves an EPC of 8.6 nJ at 0.82 V operating voltage. This is the second lowest EPC reported, for a manufactured ASIC that operates on MNIST. As the ConvCoTM ASIC is fully digital, the classification rate and thereby also the peak power consumption, can be scaled by adjusting the clock frequency. At a clock frequency of 1 MHz, the power consumption is 81 $\mu$W.
@@
Most ULP accelerators are based on analog techniques [20, 21, 9, 5]. The test accuracy of such solutions is susceptible to variations in process parameters, supply voltage and operating temperature (PVT). All-digital solutions, like our ConvCoTM ASIC and [6, 7, 8], avoid such non-ideal effects. In addition, these accelerators are easier to integrate in larger SoCs, and can more easily be scaled and re-targeted to other technologies. This applies for CNNs as well as TMs.
@@
In the ConvCoTM accelerator, the combinational clause logic draws only a small amount of energy compared to the clock tree of the inference-core DFFs. This is the main reason why the CSRF technique (see Subsection IV-D) did not provide a greater reduction in EPC. Other convolutional TM architectures, that operate at higher clock frequencies, process more literals in parallel, and utilize sparsity in TM models, can potentially benefit more from the CSRF technique.
VII Scaling
The ConvCoTM accelerator’s architecture described in this paper, is highly suited for scaling, and can be modified to operate on larger and more complex images. There is active research on how to apply the TM algorithm for larger and more complex images. The TM version that currently achieves the best test accuracy on CIFAR-10 is TM Composites [17, 18]. Here, several TM Specialists are applied to an image sample. Examples of specializations are different booleanization techniques and convolution window sizes. After an image has been processed, the class sums from each TM Specialist are normalized, followed by summation per class. Finally, an argmax operator is applied on the composite class sums for output prediction.
Table III shows characteristics and performance estimates for an envisaged TM Composites ASIC design, intended for inference operation on CIFAR-10 images. It is assumed that one configurable TM module is implemented in the ASIC, and models for four different TM Specialist configurations are applied in sequence, through transferring a new model from on-chip ultra-low-power RAM per configuration.
Booleanization of the images, for the different TM Specialists, is assumed to be performed by dedicated logic in the HW accelerator. We assume a timing-optimized data interface, with line buffers [42], allowing the patch generation to start as soon as a sufficient number of image lines have been transferred to the accelerator.
For the TM Specialists, we assume an average number of literals per patch of 1000, and by utilizing the technique described in [43], the number of literals is limited to 16 per clause. For the TA actions, the model for each TM Specialist therefore needs to include 16 ten-bit literal addresses per clause.
We assume the same 65 nm CMOS technology is used as for the current ConvCoTM ASIC. The chip area for the CIFAR-10 accelerator is estimated by multiplying the chip area of the current ASIC (2.7 mm2) with the ratio, $R$, given by the model size of one TM Specialist (32.5 kilobyte) divided by the model size of the ASIC in this paper (5.6 kilobyte). We therefore get $R=32.5/5.6\approx 5.8$. This is a reasonable assumption because the model storage in registers (for the active TM Specialist) and the clause logic, dominate the chip area. We further assume 2 mm2 in additional chip area for booleanization logic, additional adders for class sum generation, and RAM modules for model storage.
We estimate that each TM Specialist will need approximately 1000 clock cycles for the processing per sample, including booleanization. In addition, before a TM Specialist can start operating, its model has to be loaded from on-chip RAM. We assume the model is organized in memory such that a transfer rate of 32 bytes per clock cycle is enabled. Thus, we will need approximately 1020 clock cycles for a complete model transfer. A single TM Specialist will therefore require in total about 2020 clock cycles to process a sample, and with four TM Specialists we will need approximately 8080 clock cycles for a complete classification. With a system clock of 27.8 MHz (as for the current ConvCoTM ASIC), the classification rate would be $808027.8\times 10^{6}/$ 8080 FPS $\approx 3440$ FPS.
An estimate of the power consumption is found by multiplying the power consumption of the current ASIC’s accelerator core, operating at 27.8 MHz (1.15 mW), with the factor $R$. During model loading and booleanization, we assume power consumption similar to that during inference. An estimate of the energy per classification is then $R\times 1.15$ mW/(3440 FPS) $\approx 1.9~{}\mu$J/frame.
TABLE III: Envisaged ConvCoTM inference ASIC Configuration and Performance for operation on the CIFAR-10 Dataset, based on the TM Composites architecture.
Parameter Value Comment
Number of TM 4 $4\times 4$ color thermometer,
specialists $3\times 3$ color thermometer,
$32\times 32$ histogram of gradients,
$10\times 10$ adaptive thresholding.
Number of clauses 1000
Number of included 16 We assume 1000 literals
literals per clause per patch, and literal limitation
based on [43].
Model size: TA actions 20 1000 clauses, each with 16
per TM Specialist kilobyte literals. The address
length per literal is 10 bits.
Model size: Weights 12.5k 10 classes, 1000 clauses
per TM Specialist kilobyte and 10-bit weights.
Complete model size 130 For four TM Specialists.
kilobyte
Classification rate 3440 FPS Sequential operation of the
TM Specialists, 27.8 MHz
system clock, complete model
stored on-chip, without
system processor overhead.
Test accuracy 79% Estimate for CIFAR-10.
Accelerator core 17.7 mm2 65 nm CMOS.
Accelerator core 6.7 mW Inference mode,
power consumption 27.8 MHz system clock.
Accelerator core 1.9 $\mu$J Inference mode,
energy per 27.8 MHz system clock.
classification.
VIII Conclusion